# Learning relevant contextual variables within Bayesian optimization

### v3

@@ D.2 Synthetic test functions
defined over $\mathcal{V}=[0,1]^{6}$. The second, fifth, and sixth variables were considered as design variables, while the first, third, and fourth variables were considered as contextual variables. 6 noise variables were added. Table S1 provides the results of a Sobol global sensitivity analysis performed using evaluations of the function collected over a grid of $N=917504$ samples (Sobol, 2001). Adding up the first order indices for design and contextual variables separately leads to $S_{\mathbf{x}}\approx 0.124$ and $S_{\mathbf{z}}\approx 0.196$. This means that with respect to first-order interactions, contextual variables have more impact than design variables, in this synthetic example. One should notice however that these indices are computed across the whole search space and not specifically at the optimum.

### v4

@@ E.2 Synthetic test functions
defined over $\mathcal{V}=[0,1]^{6}$. The second, fifth, and sixth variables were considered as design variables, while the first, third, and fourth variables were considered as contextual variables. 6 noise variables were added. Table S1 provides the results of a Sobol global sensitivity analysis performed using evaluations of the function collected over a grid of $N=917504$ samples [Sobol, 2001]. Adding up the first order indices for design and contextual variables separately leads to $S_{\mathbf{x}}\approx 0.124$ and $S_{\mathbf{z}}\approx 0.196$. This means that with respect to first-order interactions, contextual variables have more impact than design variables, in this synthetic example. One should notice however that these indices are computed across the whole search space and not specifically at the optimum.